# Generalization of Gibbs and Langevin Monte Carlo Algorithms in the Interpolation Regime

### v1

@@ 3.3 Loss functions and secondary loss functions
$\kappa\left(\frac{1}{n}\sum_{i}f(h,x_{i}),\mathbb{E}_{x}[f(h,x)]\right)\leq\frac{1}{n}\left(\Gamma(h,\mathbf{x},\bm{\beta})+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)$
and with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$
@@
$\kappa\left(\frac{1}{n}\sum_{i}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[f\left(h,x_{i}\right)\right],\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\mathbb{E}_{x}\left[f\left(h,x\right)\right]\right)\leq\frac{1}{n}\left(\Gamma\left(h,\mathbf{x},\bm{\beta}\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right).$
For the second part, we used the joint convexity of $\kappa$. Under the conditions of this corollary, the second inequality becomes
@@
$\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\mathbb{E}_{x}\left[f\left(h,x\right)\right]\leq\kappa^{-1}\left(\frac{1}{n}\sum_{i}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[f\left(h,x_{i}\right)\right],\frac{1}{n}\left(\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)\right)$
with an analogous version for the single-draw case. This is how we compute bounds in our experiments. For an illustration, please refer to C.2.1.
Here $f$ plays the role of a secondary loss function, typically different from the loss function $\ell$, which defines the Gibbs posterior. In applications one would define the Gibbs posterior in terms of a differentiable loss function $\ell$ and approximate it with a suitable Monte Carlo method. For classification, however, one is interested in bounding the 0-1 loss obtained by comparing $\ell$ to some threshold $\theta$. Then (momentarily changing notation by replacing $x\in\mathcal{X}$ by $\left(x,y\right)$, where $y$ is the label corresponding to $x$) one can define $f\left(h,\left(x,y\right)\right)=1_{\left(-\infty,\theta\right)}\left(y\ell\left(h,x\right)\right)$ to obtain a bound on the expected 0-1 loss.
4 Bounds for Langevin Monte Carlo
For this section, we assume $\mathcal{H}=\mathbb{R}^{d}$ and an isotropic Gaussian prior $\pi$ of width $\sigma$. We condition on the training data $\mathbf{x}$, reference to which we omit. The Gibbs posterior is an idealization, from which it is impossible to sample directly. Nevertheless a number of works (Raginsky et al., 2017; Dalalyan & Karagulyan, 2017; Brosse et al., 2018; Vempala & Wibisono, 2019; Dwivedi et al., 2019; Nemeth & Fearnhead, 2021; Balasubramanian et al., 2022) discuss algorithms (SGLD, ULA, MALA, etc, here summarized as Langevin Monte Carlo (LMC)), capable of approximating a probability measure $\nu$ on $\mathbb{R}^{d}$ of the form $\nu\propto\exp\left(-V\right)$ or some nearby limiting distribution. In the following, we discuss one of these algorithms.
4.1 ULA
@@
We focus on the results of Vempala & Wibisono (2019), which do not require convexity of $V$ and instead assume that the measure $\nu$ satisfies a log-Sobolev inequality (LSI) in the sense that for all smooth $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$
$\mathbb{E}_{h\sim\nu}\left[f^{2}\left(h\right)\ln f^{2}\left(h\right)\right]-\mathbb{E}_{h\sim\nu}\left[f^{2}\left(h\right)\right]\ln\mathbb{E}_{h\sim\nu}\left[f^{2}\left(h\right)\right]\leq\frac{2}{\alpha}\mathbb{E}_{h\sim\nu}\left[\left\|\left(\nabla f\right)\left(h\right)\right\|^{2}\right]$ (5)
@@
for some $\alpha>0$. An LSI is satisfied when $V$ is strongly convex, but, importantly, also for measures which are bounded perturbations of measures satisfying an LSI (Holley & Stroock (1986)). Vempala & Wibisono (2019) give further examples and a list of references for measures, which are not log-concave and satisfy an LSI. Raginsky et al. (2017) show, that under dissipativity conditions of the loss the Gibbs posterior $G_{\beta}\left(\mathbf{x}\right)$ satisfies an LSI with constant independent of $\mathbf{x}$.
Consider the iterative algorithm
@@
where $\epsilon$ is a step size, the $\xi_{t}\sim\mathcal{N}\left(0,I\right)$ are independent Gaussian vectors and $h_{0}$ is drawn from some initial distribution $\nu_{0}$. Some authors call this algorithm simply LMC, for Langevin Monte Carlo. We call it ULA, alongside Durmus & Moulines (2017),Dwivedi et al. (2019) and Vempala & Wibisono (2019), for Un-adjusted Langevin Algorithm, because it misses the Metropolis-type accept-reject step, which would guarantee that the invariant distribution is indeed the Gibbs posterior. A popular variant of ULA is Stochastic Gradient Langevin Dynamics (SGLD) (Welling & Teh, 2011; Raginsky et al., 2017) where the gradient is replaced by an unbiased estimate, typically realized with random minibatches. Here, we restrict ourselves to ULA with a constant step size, because it has the least number of parameters to adjust, but in experiments we also use the more computationally efficient SGLD.
As $\epsilon\,{\rightarrow}\,0$, ULA recovers the Continuous Langevin Dynamics (CLD) given by the stochastic differential equation
$dh_{t}=-V\left(h_{t}\right)dt+\sqrt{2}dB_{t},$
where $B_{t}$ is centered standard Brownian motion in $\mathbb{R}^{d}$. CLD converges exponentially to the Gibbs posterior (Chiang et al., 1987). For $\epsilon>0$, the distribution $\nu_{\epsilon,t}$ of ULA converges as $t\rightarrow\infty$ to a biased limiting distribution $\nu_{\epsilon}$ which is generally different from $\nu$, but expected to be closer to $\nu$ as $\epsilon$ becomes smaller. Vempala & Wibisono (2019) use the LSI assumption and coupling to control the difference between CLD and ULA along their path and prove the following result.
Theorem 4.1.
@@
$KL\left(\nu,\nu_{\epsilon,t}\right)\leq e^{-\alpha\epsilon t}KL\left(\nu,\nu_{0}\right)+\frac{8\epsilon dL^{2}}{\alpha}.$
The first exponential term is due to the mismatch of the initial distribution and $\nu$. The second term bounds the divergence between the limiting distribution $\nu_{\epsilon}\,\$and $\nu$. Similar results exist under different conditions on the potential $V$. Cheng et al. (2018) for example require $V$ to be strongly convex outside of a ball instead of the log-Sobolev inequality and gives bounds in terms of the $W_{1}$-Wasserstein metric. Raginsky et al. (2017) give bounds for $W_{2}$ under dissipativity assumptions. The next corollary adapts Theorem 4.1 to the situation studied in this paper.
Corollary 4.2.
@@
where $h_{0}\sim\nu_{0}$ and the $\xi_{t}\sim\mathcal{N}\left(0,I\right)$ are independent Gaussian random variables. Let $D\left(\beta\right)=KL\left(G_{\beta},\nu_{0}\right)$ and let $\nu_{\beta,\eta,t}$ be the distribution of $h_{t}$ after $t$ steps. Then,
(i) $KL\left(G_{\beta},\nu_{\beta,\eta,t}\right)\leq e^{-\alpha\eta t/\beta}D\left(\beta\right)+\frac{8\eta d}{\beta\alpha}\left(\beta R+\frac{1}{\sigma^{2}}\right)^{2}.$
(ii) $W_{2}\left(G_{\beta},\nu_{\beta,\eta,t}\right)\leq\frac{2}{\alpha}e^{-\alpha\eta t/\beta}D\left(\beta\right)+\frac{16\eta d}{\beta\alpha^{2}}\left(\beta R+\frac{1}{\sigma^{2}}\right)^{2}.$
(iii) $d_{TV}\left(G_{\beta},\nu_{\beta,\eta,t}\right)\leq e^{-\alpha\eta t/\left(2\beta\right)}\sqrt{D\left(\beta\right)}+2\sqrt{\frac{\eta d}{\beta\alpha}}\left(\beta R+\frac{1}{\sigma^{2}}\right).$
@@
We make the identifications $V\left(h\right)=\beta\hat{L}\left(h\right)+\left\|h\right\|^{2}/\left(2\sigma^{2}\right)$, $\epsilon=\eta/\beta$ and $L=\beta R+\frac{1}{\sigma^{2}}$. Then, $\nu=G_{\beta}$ with Gaussian prior of width $\sigma$, ULA becomes (7) and (i) follows directly from Theorem 4.1, (ii) follows from Otto & Villani (2000) and the LSI assumption, and (iii) follows from Pinsker’s inequality (see e.g. Boucheron et al. (2013)). ∎
4.2 Stability of the bounds
We now show the stability of our bounds for approximation in total variation and $W_{2}$-Wasserstein metrics, under boundedness or Lipschitz conditions. Together with Corollary 4.2 this implies bounds for the algorithm defined in (7).
We assume that there is a target approximation $\nu_{\beta}\left(\mathbf{x}\right)$ of $G_{\beta}\left(\mathbf{x}\right)$, for which we want to compute a high probability bound, either for the single draw version on $F\left(h,\mathbf{x}\right)$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$, or, for the classical PAC-Bayesian version, on $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ as $\mathbf{x}\sim\mu^{n}$. It is not surprising that the single-draw bound will require a much closer approximation of the Gibbs posterior.
Since the bounding functional $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ depends on the Gibbs posteriors $G_{\beta_{k}}\left(\mathbf{x}\right)$ for $k\in\{1,\dots,K-1\}$, we require corresponding approximations $\nu_{\beta_{k}}$ of $G_{\beta_{k}}\left(\mathbf{x}\right)$ to compute the bound. To streamline notation, we define
$\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)=-\beta\hat{L}\left(h,\mathbf{x}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)\mathbb{E}_{g\sim\nu_{\beta_{k-1}}\left(\mathbf{x}\right)}\big[\hat{L}(g,\mathbf{x})\big]$
for $0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta$ and $\left(\nu_{0}\left(\mathbf{x}\right),\nu_{\beta_{1}}\left(\mathbf{x}\right),\cdots,\nu_{\beta_{k-1}}\left(\mathbf{x}\right)\right)\in\mathcal{P}\left(\mathcal{H}\right)^{K}$. Like $\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)$, the functional $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)$ depends on the training data, but it can also be computed by repeated execution of the algorithm (7). The next theorem states the obtained bound in terms of the approximation errors in total variation.
Theorem 4.3.
Suppose that $\mathcal{H}=\mathbb{R}^{d}$ and that there are numbers $m,M<\infty$ such that for every $\mathbf{x}$ in $\mathcal{X}^{n}$ and $h\in\mathcal{H}$ we have $\left|\ell\left(h,\mathbf{x}\right)\right|\leq m$ and $\left|F\left(h,\mathbf{x}\right)\right|\leq M$. Let $0=\beta_{0}<\beta_{1}<\dots<\beta_{K}=\beta$ and $\nu_{\beta_{k}}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ be such that $d_{TV}\left(\nu_{\beta_{k}}\left(\mathbf{x}\right),G_{\beta_{k}}\left(\mathbf{x}\right)\right)=\epsilon_{\beta_{k}}$. Then,
(i) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$

### v2

@@ B.2 Concrete PAC-Bayesian Bounds and the 0-1 Loss
$\kappa\left(\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right],\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[L\left(h\right)\right]\right)\leq\frac{1}{n}\left(KL\left(\nu\left(\mathbf{x}\right),\pi\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)$ (11)
and the corresponding inequality obtained from the inversion rule.
The inversion rule produces directly interpretable bounds, but stronger is the direct inversion using the function $\kappa^{1}:\left[0,1\right]\times\left[0,\infty\right)$ by
In the definition of $F$, using the function $\kappa$, we can use other loss functions, possibly different from the loss function $\ell$ which define the Gibbs posterior. If these loss functions satisfy the conditions of Theorem 1 in Maurer (2004), we obtain analogous bounds. In particular for binary classification we can use the 01 loss. Momentarily changing notation by replacing $x\in\mathcal{X}$ by $\left(x,y\right)$, where $y$ is the label corresponding to $x$, the 01 loss is defined as
$\hat{L}_{\textbf{01}}(h,\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}1_{\left(-\infty,0\right)}\left(y_{i}\ell\left(h,x_{i}\right)\right)$
and $L_{\textbf{01}}(h)=\mathbb{E}_{x\sim\mu^{n}}\left[\hat{L}_{\textbf{01}}(h,\mathbf{x})\right]$. We then obtain the bounds
@@
$L_{\textbf{01}}\left(h\right)$ $\leq$ $\kappa^{-1}\left(\hat{L}_{\textbf{01}}(h,\mathbf{x}),\frac{1}{n}\left(\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)\right)$
@@
$\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[L_{\textbf{01}}\left(h\right)\right]$ $\leq$ $\kappa^{-1}\left(\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}_{\textbf{01}}(h,\mathbf{x})\right],\frac{1}{n}\left(KL\left(\nu\left(\mathbf{x}\right),\pi\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)\right).$ (12)
This is the bound used in our experiments.
Appendix C Supplementary material for Section 4
C.1 Examples of LMC: CLD and ULA
@@
A number of recent works give convergence guarantees for LMC algorithms and processes (Raginsky et al., 2017; Dalalyan and Karagulyan, 2017; Brosse et al., 2018; Vempala and Wibisono, 2019; Dwivedi et al., 2019; Nemeth and Fearnhead, 2021; Balasubramanian et al., 2022; Chen et al., 2022). Here we focus on the results of Vempala and Wibisono (2019), which do not require convexity of $V$ and instead assume that the measure $\nu$ satisfies a log-Sobolev inequality (LSI) in the sense that for all smooth $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$
$\begin{split}\mathbb{E}_{h\sim\nu}\big[f^{2}(h)\ln f^{2}(h)\big]-\mathbb{E}_{h\sim\nu}\big[f^{2}(h)\big]\ln\mathbb{E}_{h\sim\nu}\big[f^{2}(h)\big]\leq\frac{2}{\alpha}\mathbb{E}_{h\sim\nu}\big[\|\nabla f(h)\|^{2}\big]\end{split}$ (13)
@@
for some $\alpha>0$. An LSI is satisfied when $V$ is strongly convex, but, importantly, also for measures which are bounded perturbations of measures satisfying an LSI (Holley and Stroock, 1986). In our applications this will be ensured for bounded losses, because of the Gaussian prior, but $\alpha$ will deteriorate as $\beta$ increases. Vempala and Wibisono (2019) give further examples and a list of references for measures which are not log-concave and satisfy an LSI. Raginsky et al. (2017) show that under dissipativity conditions of the loss the Gibbs posterior $G_{\beta}\left(\mathbf{x}\right)$ satisfies an LSI with constant independent of $\mathbf{x}$. There are a number of more recent works on this topic, in particular proximal methods (see, e.g., Chen et al. (2022)), but the work of Vempala and Wibisono (2019) is convenient for our purposes, because they guarantee convergence in relative entropy, and our emphasis is not on sampling.
Continuous Langevin Dynamics (CLD) is specified by the stochastic differential equation
$dh_{t}=-\nabla V\left(h_{t}\right)dt+\sqrt{2}dB_{t},$
where $B_{t}$ is centered standard Brownian motion in $\mathbb{R}^{d}$. The distribution of CLD converges exponentially to the Gibbs posterior under mild conditions (Chiang et al., 1987). In Section C.2, we give a convergence result for CLD adapted to temperature dependence and prior, on the condition of a log-Sobolev inequality, with convergence in relative entropy.
The Euler-discretization of CLD is the iterative algorithm
@@
where $\epsilon>0$ is a step size, the $\xi_{t}\sim\mathcal{N}\left(0,I\right)$ are independent Gaussian vectors and $h_{0}$ is drawn from some initial distribution $\nu_{0}$. Some authors call this algorithm simply LMC, for Langevin Monte Carlo. We call it ULA, alongside Durmus and Moulines (2017); Dwivedi et al. (2019); Vempala and Wibisono (2019), for Unadjusted Langevin Algorithm. A popular variant of ULA is Stochastic Gradient Langevin Dynamics (SGLD) (Welling and Teh, 2011; Raginsky et al., 2017) where the gradient is replaced by an unbiased estimate, typically realized with random minibatches. Here, we restrict ourselves to ULA with a constant step size, because it has the fewest parameters to adjust, but in experiments we also use the computationally more efficient SGLD.
The distribution $\nu_{\epsilon,t}$ of ULA converges as $t\rightarrow\infty$ to a biased limiting distribution $\nu_{\epsilon}$, which is generally different from $\nu$, but expected to be closer to $\nu$ as $\epsilon$ becomes smaller. Vempala and Wibisono (2019) use the LSI assumption to control the difference between CLD and ULA along their path and prove the following result.
Theorem C.1.
@@
$\text{KL}\left(\nu_{\epsilon,t},\nu\right)\leq e^{-\alpha\epsilon t}\text{KL}\left(\nu_{0},\nu\right)+\frac{8\epsilon dL^{2}}{\alpha}.$
The first exponential term is due to the mismatch of the initial distribution and $\nu$. Vempala and Wibisono (2019) show that $\nu_{0}$ may be chosen to make $\text{KL}\left(\nu_{0},\nu\right)$ of order $d$. The second term bounds the divergence between the limiting distribution $\nu_{\epsilon}\,\$and $\nu$. Similar results exist under different conditions on the potential $V$; Cheng et al. (2018) for example, require $V$ to be strongly convex outside of a ball instead of the log-Sobolev inequality and give bounds in terms of the $W_{1}$-Wasserstein metric. Raginsky et al. (2017) give bounds for $W_{2}$ under dissipativity assumptions. We are not aware of similar bounds for the Rényi-infinity divergence.
The next corollary adapts Theorem C.1 to the situation studied in this paper.
Corollary C.2.
@@
where $h_{0}\sim\nu_{0}$ and the $\xi_{t}\sim\mathcal{N}\left(0,I\right)$ are independent Gaussian random variables. Let $D\left(\beta\right)=\text{KL}\left(\nu_{0},G_{\beta}\right)$ and let $\nu_{\beta,\eta,t}$ be the distribution of $h_{t}$ after $t$ steps. Then
$\text{KL}(\nu_{\beta,\eta,t}\|G_{\beta})\leq e^{-\frac{\alpha\eta t}{\beta}}D(\beta)+\tfrac{8\eta d}{\beta\alpha}\big(\beta R\,{+}\tfrac{1}{\sigma^{2}}\big)^{2}$
@@
This follows directly from Theorem C.1 and the substitutions $V\left(h\right)=\beta\hat{L}\left(h\right)+\left\|h\right\|^{2}/\left(2\sigma^{2}\right)$, $\epsilon=\eta/\beta$ and $L=\beta R+\frac{1}{\sigma^{2}}$. Then $\nu=G_{\beta}$ with Gaussian prior of width $\sigma$ and ULA becomes (15). ∎
C.2 Convergence of CLD under LSI
The following is a straightforward adaptation of Theorem 1 and its proof in Vempala and Wibisono (2019). There is no claim to originality.
Lemma C.3.
Let the process $h_{\beta,t}$ on $\mathbb{R}^{d}$ be defined by the stochastic differential equation